### v1

# A combined Quantum Monte Carlo and DFT study of the strain response and magnetic properties of two-dimensional (2D) 1T-VSe2 with charge density wave

Abstract
Two-dimensional (2D) 1T-VSe2 has prompted significant interest due to the discrepancies regarding alleged ferromagnetism (FM) at room temperature, charge density wave (CDW) states and the interplay between the two. We employed a combined Diffusion Monte Carlo (DMC) and density functional theory (DFT) approach to accurately investigate the magnetic properties and response of strain of monolayer 1T-VSe2. Our calculations show the delicate competition between various phases, revealing critical insights into the relationship between their energetic and structural properties. We went on to perform Classical Monte Carlo simulations informed by our DMC and DFT results, and found the magnetic transition temperature () of the undistorted (non-CDW) FM phase to be 228 K and the distorted (CDW) phase to be 68 K. Additionally, we studied the response of biaxial strain on the energetic stability and magnetic properties of various phases of 2D 1T-VSe2 and found that small amounts of strain can enhance the , suggesting a promising route for engineering and enhancing magnetic behavior. Finally, we synthesized 1T-VSe2 and performed Raman spectroscopy measurements, which were in close agreement with our calculated results. Our work emphasizes the role of highly accurate DMC methods in advancing the understanding of monolayer 1T-VSe2 and provides a robust framework for future studies of 2D magnetic materials.
I Introduction
Strongly correlated two-dimensional (2D) magnets are a puzzling class of materials from a fundamental physics perspective. With the discovery of ferromagnetism in 2D systems such as CrI3 [1] and Cr2Ge2Te6 [2], an effort to identify and understand the underlying mechanisms of 2D magnets with a finite transition temperature () has become a highly active area of materials science. One of the most interesting and controversial 2D ferromagnetic materials isVSe2, which has a metallic 1T phase (octahedral (1T)-centered honeycombs) and a semiconducting 2H phase (trigonal prismatic (2H)-hexagonal honeycombs) [3]. Despite discrepancies of the structural properties coupled to the energetic stability (whether or not 1T vs. 2H is more favorable) [4, 5, 6, 7] that we were successfully able to resolve in our previous work using highly-accurate electronic structure methods (see Ref. 8), there are several remaining questions regarding 2D 1T-VSe2.
In a 2018 study, single layer 1T-VSe2 was synthesized on van der Waals (vdW) substrates (graphite and MoS2), where strong ferromagnetic ordering was measured above room temperature and a charge density wave (CDW) was detected with a transition temperature of 121 K [9]. This was the first study to demonstrate that the CDW of single layer 1T-VSe2 is coupled with its magnetic properties [9]. Despite room temperature ferromagnetism being reproduced in some cases, such as for a chemically exfoliated monolayer [10], there have been several conflicting reports of the magnetic properties and the interplay between the CDW state and magnetism [9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 11, 19, 20, 21, 22]. For example, some studies have found the nonmagnetic CDW phase to be experimentally favorable, with an inherent absence of ferromagnetism [13, 14, 15, 16].
Several theoretical studies have attempted to explain why these discrepancies in magnetic properties might occur, citing strain, vacancies, substrate choice, doping and chemical functionalization as possible reasons [23, 24, 25, 26, 27, 28, 29, 30, 31]. These studies also highlight the strong sensitivity of the magnetic properties of 2D 1T-VSe2 to extrinsic factors, which can provide a viable route to tune such properties. In our previous density functional theory (DFT) work, we explored the competing magnetic and nonmagnetic states in single layer 1T-VSe2 with and without charge density wave [26]. We found that there is strong competition between nonmagnetic and magnetic states in the CDW structures (with respect to the undistorted structure), with relative energies being on the the scale of one meV per formula unit (f.u.) [26]. This implies that it is possible for antiferromagnetic ordering to compete with ferromagnetic ordering and theCDW state [26].
Although our previous semi-local DFT calculations provide a good qualitative assessment of how different magnetic orderings can compete in the undistorted and CDW structures, a more quantitative answer is required. This is required in order to accurately understand the magnetic and CDW transitions in monolayer 1T-VSe2 and estimate quantities such as transition temperatures. For this reason, high-fidelity many-body techniques such as Diffusion Monte Carlo (DMC) [32] can be utilized to accurately describe the electronic and magnetic properties of 2D 1T-VSe2. DMC is a correlated electronic structure method that has a reduced sensitivity on approximations such as the exchange-correlation functional and the Hubbard [33] correction. In addition, DMC has successfully been applied to several 2D and quasi-2D systems [8, 34, 35, 36, 37, 38, 39, 40, 41, 42, 43, 44, 45, 46, 47, 48, 49]. In this work, we investigate the magnetic properties of monolayer 1T-VSe2 in various magnetic states with and without CDW through the lens of DMC and provide a thorough benchmark using several DFT functionals. In addition to providing accurate quantitative estimates of energy differences between different phases, we coupled our DMC results with classical Monte Carlo simulations to estimate magnetic transition temperatures. Finally, we used our DMC insights to perform DFT calculations of the Raman modes and compared to our own experimental results on a synthesized 1T-VSe2 flake.
II Methodology
Benchmarking DFT calculations for monolayer 1T-VSe2 were performed with the Vienna Ab initio Simulation Package (VASP) code, using projector augmented wave (PAW) pseudopotentials [50, 51]. For testing purposes, we employed a variety of exchange-correlation functionals including the local density approximation (LDA)[52, 53], Perdew-Burke-Ernzerhof (PBE)[54], the strongly constrained and appropriately normed (SCAN)[55] meta-GGA functional and the r2SCAN [56] meta-GGA functional with, and without the Hubbard correction [33], which was used to treat the on-site Coulomb interaction of orbitals of V. A k-grid of 24x24x1 was used and scaled proportionally to the supercell size, a plane-wave kinetic energy of 400 eV (increased to 800 eV for SCAN and r2SCAN) was used and at least 20 Å of vacuum was added between periodic layers. Our density functional perturbation theory (DFPT) phonon calculations to obtain the peak positions of Raman modes were also performed with VASP along with the phonopy [57, 58] package.
Our Quantum Monte Carlo (QMC) calculations for 1T-VSe2 used the exact same settings as our previous work on 1T- and 2H-VSe2 (see Ref. 8 for more specific details regarding convergence testing). DFT-PBE was used to generate the trial wavefunction (nodal surface) for DMC using the Quantum Espresso (QE) [59] code with a correction of 2 eV (PBE+ = 2 eV resulted in a trial wavefunction with the lowest total energy [8]). We used a k-point grid of 12x12x1 and a kinetic energy cutoff of 4,080 eV (300 Ry) to generate the trial wavefunction. In terms of pseudopotentials, RRKJ (OPT) potentials were used for V [60] and Burkatzki-Fillipi-Dolg (BFD) were used for Se [61]. Variational Monte Carlo (VMC) and DMC [32] simulations were conducted using the QMCPACK [62, 63] code and the Nexus [64] workflow manager was used to automate the DFT-VMC-DMC calculations. Up to two-body Jastrow parameters [65, 66, 67] were optimized using the linear method [68, 69] to minimize the variance and energy in VMC, with the goal of modeling electron correlation and reducing uncertainty in the DMC results [69, 70]. In order to compute the nonlocal part of the pseudopotentials, the locality approximation [70] was used in DMC. An optimal timestep of 0.01 Ha-1 (0.27 eV-1) [8] was used in DMC and to reduce finite-size errors, twist averaging and extrapolating to larger supercells (up to 72 atoms) was performed. A more detailed explanation of the theory behind VMC and DMC can be found in Ref. 32.
Energies of different spin configurations derived from first-principles calculations were mapped onto a classical spin Hamiltonian, characterizing three-dimensional spins arranged on a two-dimensional lattice at sites of V atom. Subsequently, this Hamiltonian served as the basis for executing classical Monte Carlo simulations aimed at estimating the transition transition temperature () for both the undistorted and CDW phases. The Hamiltonian, which has previously been used for other 2D magnetic systems [71, 72, 37], includes the bilinear isotropic exchange (), anisotropic exchange (), and easy axis single ion anisotropy (),
The indices and iterate over all magnetic sites and their corresponding first nearest neighbor (NN) magnetic sites, respectively. Only the first NN interactions are considered due to the strong localized magnetic moment on V. We have calculated for the CDW phase to be , supporting our assumptions. According to the adopted sign convention, favors isotropic spin alignment, leading to a preferred ferromagnetic (FM) over an antiferromagnetic (AFM) phase. Similarly, favors alignment of the spin -components, and favors the out-of-plane direction as the easy axis. We employ a supercell, equivalent to a nearly nm2 square cell, for the MC simulations. The spin configuration space is discretized into uniformly distributed points on the surface of a unit sphere with a fine resolution of degrees, facilitating detailed exploration of spin orientations. The MC simulation consists of numerous sweeps across all magnetic sites. During each sweep, the spin orientation at each site is updated randomly, one site at a time, and the new configuration is then either accepted or rejected based on the Metropolis algorithm. The convergence of both energy and magnetization is meticulously monitored across varying temperatures. Upon achieving equilibration, an ensemble of configurations is collected to calculate the magnetization and specific heat. We use an ensemble size of k for the CDW simulation and k for the undistorted simulation. Additional methodology details can be found in the SI.
III Results and Discussion
In this work, we specifically focused on monolayer 1T-VSe2 in its freestanding form. Motivated by the experimental discrepancies of controversial room temperature ferromagnetism and competing magnetism and CDW states in the 1T phase, we decided to perform in-depth benchmarking at the DFT level using a variety of approximations. We focused on the normal (undistorted) structures of 2D 1T-VSe2 in its FM, AFM (spins are anti-aligned in a stripy pattern) and nonmagnetic (NM) orientations. In addition to the normal unperturbed crystal structure of 1T-VSe2, we studied the distorted xx1 supercell which is a signature of the material being in its CDW state. This distorted xx1 CDW structure has previously been studied extensively with DFT and experimentally verified. Building on the work of Ref. 26, we performed a thorough DFT benchmark (using several different approaches) of the various magnetic states of the distorted CDW structure. These configurations included FM, NM, and various AFM configurations. Consistent with Ref. 26, we studied the xx1 distorted AFM-A, AFM-B1 and AFM-B2 configurations (see Fig. 1). It is important to note that the distorted AFM-A configuration is equivalent to the AFM (stripy) configuration for the undistorted structure. A summary of these different structures and magnetic configurations are given in Fig. 1.
In our previous work (Ref. 8), we determined the optimal geometry (including lattice parameter and bond distance) and relative phase stability of 2D 1T- and 2H-VSe2 in the FM orientation with DMC and benchmarked with several DFT functionals with and without the Hubbard correction. Fig. 2a) depicts the DFT benchmarking calculations for the NM, FM, and AFM orientations of the undistorted structure. We performed DFT calculations with LDA, PBE, SCAN, r2SCAN with values ranging from (0 to 3) eV, where we fully relaxed each structure in its respective magnetic orientation. As seen in Fig. 2a), which plots the energy differences with respect to the NM state for each functional, the discrepancies with each DFT method are enormous. Although qualitatively the results are somewhat similar (i.e., the FM state being lowest in energy across the board for most DFT functionals), quantitatively, the results vary drastically. In fact, the energy per formula unit (f.u.) between the NM orientation and the FM and AFM orientations can vary up to 0.4 eV. A quantitative energy difference between magnetic states is absolutely necessary to obtain magnetic exchange parameters () and therefore transition temperature () with reasonable accuracy.
Energy Difference | DMC | SCAN | r2SCAN |
(eV/f.u.) | |||
E - E | -0.07(1) | -0.054 | -0.044 |
E - E | -0.20(1) | -0.187 | -0.166 |
E - E | -0.13(1) | -0.133 | -0.122 |
E - E | -0.23(1) | -0.184 | -0.168 |
E - E | - | 0.010 | -0.005 |
E - E | - | -0.122 | -0.120 |
E - E | - | -0.161 | -0.148 |
E - E | - | -0.173 | -0.157 |
In order to overcome these shortcomings of local and semilocal DFT, we performed DMC calculations to obtain accurate total energies of the NM, FM and AFM orientations of the normal undistorted structure of monolayer 1T-VSe2. As a starting geometry for our DMC calculations, we used the structure obtained from our previous work in Ref. 8. We acknowledge that the local magnetic ordering can impact the structural properties, but for our DMC calculations we used the same structure for NM, FM and AFM calculations. In order to study the influence of using the same geometry for NM, AFM and FM on the energetics, we performed DFT calculations with the fixed FM geometry for all magnetic orientations and found a maximum energy difference of 10 meV between the fixed and relaxed structures, which is within the error bar of our DMC calculations. To reduce finite-size errors, we performed DMC calculations for two reasonably sized supercells for NM, FM and AFM and extrapolated to the thermodynamic limit. We used the following supercell sizes: 36 and 72 atoms for NM and AFM, 27 and 48 atoms for FM.The DMC energy differences and uncertainties are depicted in Fig. 2 and Table 1 (in shorthand notation). From the figure and table, we see that SCAN and r2SCAN (with no Hubbard correction) are in closest agreement with our DMC benchmark. For the FM configuration, we also find that SCAN and r2SCAN successfully reproduce the DMC geometry that was previously obtained in Ref. 8.
Although other semilocal DFT functionals such as PBE+ ( = 1, 2 eV) can do a sufficient job at reproducing our DMC geometry, they fail in terms of correctly capturing the quantitative energy differences of various magnetic states. Meta-GGA functionals (SCAN, r2SCAN) can be quite successful for 2D magnetic materials such as VSe2. SCAN and r2SCAN include the kinetic energy density (in addition to the electron density and its gradient) in the exchange-correlation functional, which allows them to capture complicated electronic interactions. These functionals can reduce the self-interaction errors of typical GGAs and improve the description of correlation effects. SCAN satisfies all 17 exact known constraints that a meta-GGA can satisfy, while r2SCAN breaks some of these exact constraints in favor of numerical stability (without significantly hindering accuracy). Although these functionals can simultaneously capture multiple properties such as energy differences between magnetic phases and optimal geometry, SCAN and r2SCAN have been widely reported to overestimate on-site magnetic moments [73, 74, 75, 76]. In the case of 1T-VSe2, we found the DMC on-site magnetic moment of the V atom to be 1.06(2) in Ref. 8. We found the on-site moment to be 1.40 with SCAN and 1.43 with r2SCAN. This discrepancy can be due in part to the tendency of SCAN/r2SCAN to overestimate magnetic moments [73, 74, 75, 76]. It can also be due in part to the limitations of using a single determinant wavefunction in our DMC calculations [77, 78, 79, 80]. Nonetheless, due to the success of multi-property predictions with SCAN and r2SCAN with respect to our DMC benchmark calculations, we decided to use SCAN to obtain more complex quantities (that are difficult to obtain with DMC) such as magnetic anisotropy energies. For computationally demanding simulations such as for the Raman active modes where it is not computationally tractable to use SCAN or r2SCAN, we used PBE+ ( = 1 eV) due to its relative qualitative agreement with DMC results (in terms of structural properties).
Fig. 2c) depicts the results for the CDW distorted structures with various magnetic orientations. The initial supercells for these distorted structures were obtained from Ref. 26, where the V and Se atoms were initially displaced by 0.12 and 0.18 Å respectively (similarly to Ref. 15) prior to atomic relaxation. For our DFT benchmarking calculations, we adapted a slightly different approach for the distorted CDW structures (in contrast to the undistorted cells). For these supercells, we modified the lattice constant to be a direct multiple of the unit cell geometry we obtained in our previous work (Ref. 8) and fixed the cell dimensions. We then allowed the atomic positions (with initially displaced atoms) to relax with each respective DFT functional. Similarly to the undistorted structures in Fig. 2a), there is a massive discrepancy in relative energy between each magnetic orientation (with respect to the NM undistorted cell) of the distorted CDW structures (up to 0.5 eV). We observe that the FM orientation is the lowest energy magnetic state for all functionals depicted in Fig. 2b). In contrast to the undistorted structure in Fig. 2a), the energy differences between the FM and various AFM orientations is much smaller. Our results are in good qualitative agreement with the DFT calculations performed in Ref. 26, but are shifted slightly in energy which can be attributed to fixing the lattice constant to the DMC obtained value. Similar to the undistorted case, we also find excellent qualitative agreement between SCAN/r2SCAN and PBE+ ( = 1 eV) for the energy differences between FM and AFM configurations of distorted structures. To provide more insight on these DFT calculations, we performed DMC simulations for just the FM CDW distorted structure at supercells of 30 and 60 atoms and extrapolated to the thermodynamic limit. For the starting geometry of this structure, we optimized the initially displaced atomic coordinates with r2SCAN and used a direct multiple of the unit cell geometry we obtained in our previous work (Ref. 8). This DMC energy difference and uncertainty is shown (with respect to the undistorted NM structure) in 2c) and Table 1. We find that the DMC energy difference lies between PBE+ ( = 1 eV) and PBE+ ( = 2 eV). Additionally, we find that the SCAN/r2SCAN and LDA+ ( = 3 eV) energy differences are close to the DMC value.
In addition to benchmarking the energy differences with DFT and DMC for the distorted CDW structures with various magnetic orientations, we decided to study the impact that each DFT functional has on the structural distortions (initially displaced atoms). Interestingly, we found that when the atomic coordinates of some of the distorted CDW structures (with specific DFT functionals and Hubbard corrections) are relaxed, the distortions are significantly reduced and almost disappear. This is most apparent for the distorted AFM-A orientation. For example, when SCAN, r2SCAN, PBE+ ( = 1, 2, 3 eV) and LDA+ ( = 2, 3 eV) are used to relax the AFM-A structure, the distortions disappear completely and the structure relaxes to the normal (undistorted) AFM (stripy) structure. This implies that the stripy AFM ordering cannot exist in the distorted structure. Interestingly, for all DFT functionals considered in this work, the distortions did not disappear from the NM-CDW structure In contrast, we see the distortion disappear in CDW-AFM-B1 for LDA+ and PBE+ ( = 3 eV) and in CDW-AFM-B2 for LDA+ ( = 3 eV). Regarding the CDW-FM structure, we see a significant reduction of the distortion for LDA+ ( = 0, 3 eV) and PBE+ ( = 0, 2, 3 eV). The disappearance of the CDW distortion may be due in part to performing these DFT calculations at a fixed lattice constant predicted by DMC (from Ref. 8) that is different than the lattice constant computed by each respective functional, which can induce strain. The fact that the geometry predicted by SCAN/r2SCAN and PBE+ are closest to the DMC value could be an explanation for why the distortion does not disappear for those functionals (since minimal strain is induced). From these results, we can infer that FM in the distorted-CDW structure is highly sensitive to strain (which is discussed later on in this manuscript). This highlights the competitive nature and interplay of structural parameters, strain mechanisms and magnetic ordering that can occur within the 1T-VSe2 monolayer.
For comparison sake, we analyzed the distortions in the FM-CDW structure obtained from SCAN/r2SCAN (fixed to the scaled DMC lattice constant) to the distortions obtained from PBE calculations in Ref. [26] (where the lattice constant was fixed to 3.36 Å). These comparisons are illustrated in Table S1, where we display the deviation in nearest neighbor bond distances (V-V) with respect to the undistorted bond distances (calculated with SCAN, r2SCAN and PBE respectively) for all of the V-V bonds in the honeycomb lattice (6 bond distances for 5 different V atoms in the CDW supercell). We find that the distortions are reduced in the SCAN/r2SCAN calculations when compared to the previous PBE results [26], which can again be due in part to fixing the lattice constant and inducing strain. Nonetheless, our DMC energies indicate that undistorted FM and CDW-FM are quite close in energy (-0.20(1) eV vs. -0.23(1) eV, with respect to the undistorted NM phase). We propose that the -0.20 eV is coming from the magnetic instability while the -0.03 eV is coming from the CDW, which is consistent with the fact that the magnetic moments (SCAN) between undistorted FM and CDW-FM are nearly identical.
From our DMC insight, we went on to estimate the bilinear isotropic exchange (), anisotropic exchange (), and easy axis single ion anisotropy (). Since the incorporation of spin-orbit within DMC is a relatively new development and the spin-orbit contributions for 1T-VSe2 are within the DMC uncertainty, we performed collinear DMC calculations (where spins can orient either up or down) for the undistorted FM and AFM orientations (depicted in Fig. 2 and Table 1). For the undistorted structure, we obtain a value of 71(10) meV (no spin-orbit contribution). Taking the insight from our DMC and DFT calculations, where we observed that SCAN reproduced our DMC benchmarking results, we went on to perform spin-orbit SCAN calculations to explicitly obtain anisotropy parameters and (see SI for more detailed information). From these calculations, we obtained a value of -0.77 meV and a value of 0.27 meV. With regards to the distorted-CDW structure, we performed a DMC benchmark calculation for the FM orientation (see Fig. 2 and Table 1). Due to the reasonable agreement with DMC and SCAN and for consistency with the undistorted case, we went on to calculate exchange and anisotropy parameters for the distorted structure using spin-orbit SCAN calculations. In the case of SCAN (since the AFM-A CDW-distorted structure relaxed to the undistorted AFM orientation), the AFM-B2 configuration is the ground state AFM configuration. Using our CDW-FM and CDW-AFM-B2 noncollinear SCAN calculations, we obtained a obtained a value of 21 meV, a value of -0.57 meV, and a value of -0.77 meV (see SI for more details). We went on to use these accurately computed magnetic exchange and anisotropy parameters for our spin Hamiltonian model shown in Eq. 1, where we performed classical Monte Carlo simulations to gain insight on the magnetic phase transitions of the undistorted and distorted structures.
Fig. 3 illustrates the magnetic phase transitions for both (a) undistorted and (b) charge density wave (CDW) phases. In each case, the exchange interaction is predominant over other interactions. Notably, is over three times higher in the undistorted phase, measuring 71 meV, compared to 21 meV in the CDW phase. This pronounced disparity in interaction strength indicates that spins in the undistorted structure align more strongly with neighbors, leading to an extended correlation length and larger clusters of aligned spins, requiring more thermal energy to disrupt these interactions and disorder the aligned spins. This implies that there exists two potential magnetic phase transitions for freestanding (no substrate, strain or defects) 1T-VSe2, one for the undistorted monolayer below room temperature ( 100 K less than the transition temperature reported in Ref. 9) and one for the CDW-distorted (slightly below liquid N2). These results indicate that the magnetic phase transition occurs below the CDW transition temperature, which has been reported to be above 121 K. This is expected since the CDW phase must exist before the CDW-FM phase can exist by further cooling. The small energy scale between the CDW-distorted phases and the undistorted phases (reported in our current work and previous work[26]) poses a challenge in understanding which phase is energetically more favorable. This small energy scale opens the door to tune the energetic stability of various magnetic states (with and without CDW-distortion) through external mechanisms such as strain, defects, substrate engineering or temperature.
For a deeper analysis of how structural parameters are related to the energetic favorability of different magnetic states, we performed biaxial strain calculations for the undistorted and distorted structures with the most reliable DFT functional (in our case, SCAN was most reliable for reproducing DMC geometries and quantitative DMC energy differences). These results are depicted in Fig 4. It is important to note that for the CDW distorted structures, we kept the displaced atomic positions fixed while applying biaxial strain, and excluded the AFM-A orientation since it relaxed to the undistorted AFM (stripy) orientation with SCAN. We also kept the relative atomic positions fixed for the undistorted configurations while varying the lattice.
From Fig. 4a), we observe that the CDW-distorted structure in FM and AFM configurations is more sensitive to small amounts of strain in contrast to the undistorted structure. In other words, the energy differences between the strain curves are much smaller for the CDW-distorted structures (compared to undistorted). Notably, we observe that for positive values of strain (lattice expansion), the FM in the undistorted and CDW-distorted structures becomes more stable. For the undistorted structure, we see that the energy difference between the FM state and AFM state drastically increases with the application of positive (tensile) strain. We observe that with small amounts of compressive strain, the undistorted FM structure becomes more favorable than the CDW-distorted FM structure, which can be a promising route to stabalizing the FM in the undistorted structure. For larger amounts of compressive strain ( ), AFM order begins to become more favorable for the undistorted and CDW-distorted (specifically AFMB1) structures.
To further understand this phenomena, we computed under strain for the energy differences depicted in Fig. 4a) for the undistorted and CDW-distorted structures. We then went on to estimate under strain using these values of for additional Monte Carlo simulations (similar to those reported in Fig. 3). It is important to note that for these Monte Carlo simulations under lattice strain, we did not include and in the Hamiltonian (due to increased computational demand of spin-orbit SCAN calculations). Since the driving force behind the is , we do not expect major fluctuations in from and with applied strain. These results are depicted in Fig. 4b). In addition to the strong competition between the undistorted FM and the CDW-FM structures (dotted purple and orange curves in Fig. 4a)), there is potential strain-tunability within each FM ground state (with and without CDW). In Fig. 4b), we see that can be enhanced in the CDW-FM state by small amounts of biaxial expansion or compression, with the approaching values above 120 K. With regards to the undistorted FM structure, we find that tensile strain can enhance the to be above room temperature, but there is strong competition with the FM-CDW state for small values of tensile strain, with the CDW-FM becoming significantly more energetically favorable after strain. In addition to the tunability of magnetic states, it was revealed from anharmonic phonon calculations that the CDW order can be tuned with small amounts of lattice strain (as little as 1.5 ), highlighting the strong substrate dependence [81]. For example, common substrates for VSe2 [9] such as MoS2 (a = b = 3.16 Å) and highly oriented pyrolytic graphite (a = b = 2.46 Å) have significant lattice mismatch, which can induce strain (since the lattice constant of monolayer 1T-VSe2 is estimated to be 3.41 Å[8]). As an alternative to substrate engineering, substitutional doping can be a promising route to inducing tensile strain in 1T-VSe2, which can stabalize and even enhance the magnetic properties in both the undistorted and CDW-distorted phases.
300 K | Bulk FM | ML FM | ML AFM | NM |
A1g | 199 | 197 | 197, 139 | 208 |
Eg | 137 | 135 | 130 | |
Bg | 133 | |||
CDW-ML | NM | FM | ||
A1g | 204 | 205 |
For additional benchmarking purposes, we performed DFPT calculations to obtain the assignment and peak positions of the Raman modes of 1T-VSe2 for the distorted and undistorted structures in various magnetic and nonmagnetic configurations at 300 K. Due to the higher computational cost and convergence issues of SCAN, we performed these DFPT calculations with PBE+ ( = 1), since this functional was able to match our DMC benchmark for lattice geometry. The logic behind using PBE+ ( = 1) stems from the fact that if it can correctly reproduce our lattice geometry, it can accurately capture vibrational properties. Table 2 depicts the Raman active modes calculated with DFT for the bulk and monolayer structures (with and without CDW distortions). For the undistorted structure, we performed calculations for the FM, AFM and NM orientations and for the CDW-distorted structure we considered the FM and orientations. As seen in Table 2, each structure has a pronounced A1g peak ranging from (197 to 208) cm-1. Due to a lack of symmetry in the CDW-distorted structure, we could only identify this A1g, but it is entirely possible other Raman modes can appear experimentally. For the undistorted structure (bulk and monolayer), we observe an peak ranging from (130 to 137) cm-1 for the FM and NM orientations. The most distinguishing features are apparent for the AFM orientation of the monolayer, where we observe an additional A1g peak at 139 cm-1 and a Bg peak at 133 cm-1.
To validate our DMC-informed DFT calculations of the Raman modes, we grew a VSe2 crystal flake by Chemical Vapor Transport (CVT) (see SI for details). The 1T phase of VSe2 was confirmed by X-Ray diffraction (XRD) (see Fig. S1). Thin flakes of 1T-VSe2 were exfoliated and encapsulated in h-BN (Fig. S2(a)) and subsequent Raman spectra measurements were performed. The Raman spectra were averaged (30 seconds, 20 accumulations) at four distinct positions on a 1T-VSe2 flake capped with h-BN. Fig. 5(a) shows an optical image with the approximate locations where the spectra were measured and Fig. 5(b) shows the corresponding spectra. The thinnest region measured to be about 1.7 nm by atomic force microscope (AFM) Figure S2, position 3, was used for comparison to DFT predictions. To assign the phonon modes polarized Raman spectra were collected from the 1.7 nm thick region (position 3). Data is shown in both the parallel and cross configuration, i.e., the incoming laser excitation and the Raman scatter were either parallel to one another, or at 90-degrees, confirming the respective E and A symmetry of the modes of the VSe2. In the 1.7 nm thick region we identify two modes; the A1g at 206.7 cm-1 0.4 and the Eg at 137.5 cm-1 0.3 in good agreement with the two modes predicted by DFT. The other features correspond to the Si/SiO2 substrate, as can be seen when compared to measurement of the blank substrate (position 4). However, in the bulk-like region on the h-BN capped flake, position 1, there are additional observable features to the right of the A1g mode which differ from the edge-like feature observed in position 3 coming from the substrate. Previously a broad mode at about 257 cm-1 has been attributed as an Eg mode of the 1T phase in a bulk sample [82], however, we observe in some cases two distinct modes at 235.3 cm-1 and 253.2 cm-1 in the bulk-like region in the h-BN capped flake and at 229.8 cm-1 and 250.3 cm-1 in the bulk crystal, shown in Figure S3. Since these are only observed in the bulk-like capped region/crystal, and are not predicted by the DFT results, these modes are likely due Se [83, 84, 85]. We conclude that the 1T phase of VSe2 only has two phonon modes which we identify as the higher intensity A1g and the lower intensity Eg.
IV Conclusion
We have provided a comprehensive investigation of the magnetic properties and strain response of monolayer 1T-VSe2 using DFT and highly accurate DMC methods. By extensively benchmarking various DFT functionals against our DMC results, we demonstrated the significant impact of the exchange-correlation functional on the accuracy of magnetic properties and energetic stability of various magnetic states of 2D 1T-VSe2, with and without CDW. Our high-fidelity results aim to resolve previous discrepancies that exist in the theoretical and experimental literature for 1T-VSe2. Through classical Monte Carlo simulations informed by DMC and DFT, we estimated the magnetic transition temperatures for the undistorted (228 K) and CDW-distorted (68 K) phases and our strain calculations indicated that small amounts of biaxial strain can enhance the transition temperature, which provides a viable route for engineering the magnetic properties. Furthermore, our Raman spectroscopy experiments on exfoliated flakes of 1T-VSe2 validate our predictions and provide insight on the intrinsic strain mechanisms. Our work underscores the important role of highly accurate many-body methods such as DMC in describing the electronic and magnetic properties of 2D materials. We hope our combined approach can pave the way for future explorations of correlated 2D magnetic materials.
V Notes
Please note that the use of commercial software (VASP) does not imply recommendation by the National Institute of Standards and Technology. Certain commercial equipment or materials are identified in this paper to adequately specify the experimental procedures. In no case does the identification imply recommendation or endorsement by NIST, nor does it imply that the materials or equipment identified are necessarily the best available for the purpose.
VI Competing interests
The authors declare no competing interests.
VII acknowledgments
This work was supported by the National Science Foundation through the Division of Materials Research under NSF Grant No. DMR-2213398. The authors thank the National Institute of Standards and Technology for funding, computational, and data-management resources. Work by N.G. was supported by the National Institute of Standards and Technology Summer Undergraduate Research Fellowship program. Work by K.S. and J.T.K. (discussion, analysis of QMC calculations) was supported by the U.S. Department of Energy, Office of Science, Basic Energy Sciences, Materials Sciences and Engineering Division, as part of the Computational Materials Sciences Program and Center for Predictive Simulation of Functional Materials. Work by T. B. (discussion, analysis of DFT+U calculations) was supported by the U.S. Department of Energy, Office of Science, Basic Energy Sciences, Materials Sciences and Engineering Division. F.A. and B.D. would like to acknowledge the support of Mr. Peter Morrison and the Office of Naval Research under Contract No. N00014-21-WX-01248. Research at the United States Naval Academy (M.E.J.) was supported by the Office of Naval Research under Contract No. N0001423WX02132.